OpenStudy (anonymous):

Mulitply and Simplify ( 9 + 2v3 ) ( 9 - 2v3 ) v stands for sqrt A.69 - 4v3 B. 81 - 4v3 C. 69

OpenStudy (anonymous):

I think its B.

OpenStudy (anonymous):

did you actually foil or just guess?

OpenStudy (anonymous):

i distributed

OpenStudy (anonymous):

it is similar to (a+b)(a-b)=a^2 -b^2 ( 9 + 2v3 ) ( 9 - 2v3 ) = 9^2 - (2v3)^2 = 81 - 12=69

OpenStudy (anonymous):

show how you distributed for your next question so your work can be checked